If a 165-mile trip took 7 1/2 gallons, of gas, how many miles can be driven with a full tank of 16 gallons?
Answer by rfer(16322)

165/7.5=x/16
7.5x=165*16
7.5x=2640
x=2640.7.5
x=352 mi